OpenStudy (anonymous):

An antiderivative of 3x^2/(2x^3-5) is:

OpenStudy (hba):

take its integral

OpenStudy (slaaibak):

\[3x^2 \over 2x^3 - 5\] ?

OpenStudy (slaaibak):

Integrate by substitution.

OpenStudy (anonymous):

is look a quotant rule to me

OpenStudy (slaaibak):

no. it's integration, not differentiation

OpenStudy (campbell_st):

this is a log function let u = 2x^3 - 5 du/dx = 6x^2 1/2 du = 3x^2 so you are finding the antiderivative of \[\frac{1}{2} \int\limits \frac{du}{u} = \frac{1}{2} \ln( u) + c\]